Find an equation of the line that passes through left parenthesis 2 comma 5 right parenthesis(2,5) and is perpendicular to y equals 2 x minus 7y=2x−7.
Answer by ewatrrr(24785)

y = 2x - 7 Green Line
m = 2
New Line perpendicular to it m = -1/2 (negative reciprocal of m = 2)
New Line passes through (2,5)
y - 5 = (-1/2)(x-2)
y = (-1/2)x + 6 Blue Line